# Reduction-induced Variation of Partial Von Neumann Entropy

### v2

@@ Abstract
The organization and structure of bipartite mixed-state quantum entanglement (QE) is more complicated and less well understood than that of bipartite pure-state QE. Bipartite mixed-state QEs and their measures play a crucial role in both theory and practical applications. Existing measures comprise two classes: one class involves quantifying the minimum QE and reflect the inherently complex nature of their computation, while the other is only applicable to highly limited-dimensional quantum systems. In this context, we propose a method termed Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE) to quantify QE in any bipartite states especially for any bipartite mixed-states, which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to any bipartite states. This method exhibits minimal computational complexity and broad applicability. Its intuitive and clear physical representation, combined with easy computation and wide applicability, facilitates exploring its potential applications. Furthermore, we present examples to demonstrate the superiorities of this method in identifying bipartite QE by comparing with other existing bipartite mixed-state QE measures through both their physical implications and mathematical structures.
Keywords:Quantum Entanglement (QE); Bipartite Mixed-state QE; Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE)
@@ I Introduction
The type or degree of QEs are often directly related to their revolutionary applications including quantum cryptography kk ; ng , quantum teleportation ch ; rc ; nr , quantum superdense coding nr ; sj , measurement-based quantum computation rr ; fa and quantum enhanced sensing tc ; rcp ; lp . Therefore, a deeper understanding and accurate metric of QE have become important prerequisites for the development of quantum information science and quantum technology.
@@
Nonseparable bipartite systems exhibit QE epr ; sd , which can be quantified using various measures. These include the Partial Von Neumann Entropy ch1 ; ch2 , applicable only to bipartite pure states, and Concurrence, which is confined to $2\times 2$ quantum systems wk ; au ; ka ; mf ; xjw , and Negativity, which is applicable only to $2\times 2$ and $2\times 3$ quantum systems K.Z ; gv . Additionally, Entanglement of Formation ch1 ; ch2 and Relative Entropy of Entanglement vv1 ; vv2 for bipartite quantum states, involve quantifying the minimum QE, and their assessments generally entail considerable computational burdens, especially for bipartite mixed-state QE. In general, QEs in pure states are more intuitive and easier to comprehend, while those in mixed states are more complicated and less well understood than those of pure-state QEs. Quantifying QEs in bipartite mixed states is more complex and challenging due to the contributions from classical probability distributions, making the definition and computation of QEs more intricate.
@@
Mixed states are more reflective of real-world scenarios, as factors such as environmental conditions, quantum measurements, and thermal influences can lead quantum systems to exist in mixed states. Consequently, the measures of QEs in mixed states are often more directly applicable to complex systems, aligning better with experimental conditions and practical applications. Hence, in practical research and applications, bipartite mixed-state QE and its measures hold significant practical importance, besides the theoretical significance of their underlying structure of quantum mechanics.
@@
In this context, our research aims to explore new metrics for bipartite mixed-state QE. We propose a method termed Reduction-induced Variation of Partial Von Neumann Entropy to quantify QE in any bipartite states, which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to any bipartite states. The computational complexity associated with this method is considerably low, making it highly efficient. Moreover, it has a wide range of applicability. The intuitive and clear physical representation, combined with simple computation and wide range of applications, facilitates the exploration of its specific potential applications. Subsequently, we provide examples to demonstrate the superiorities of this method in bipartite QE identification through a comparison with other established bipartite mixed-state QE measures including Partial Von Neumann Entropy, Concurrence, Negativity, Entanglement of Formation, and Relative Entropy of Entanglement, considering both their physical implications and mathematical structures.
II existing five kinds of Bipartite QE measures
@@ II.1 Partial Von Neumann Entropy for Bipartite Pure-state QE
where $\rho_{i}(i=1,2)$ represents its subsystem reduced density matrix. This bipartite QE measure, the Partial Von Neumann Entropy (PVNE) can effectively measure the degree of QE between bipartite pure states, while for mixed-state QE, it is ineffective.
@@ II.2 Entanglement of Formation for Bipartite QE and Concurrence for two-qubit QE
For bipartite QE especially for bipartite mixed-state QE, an extensively acknowledged type of measure, Entanglement of Formation (EOF) ch1 ; ch2 was introduced and is defined as
@@
over all possible ensemble realizations of the bipartite density matrix $\rho=\sum_{j}p_{j}\mid\Psi_{j}\rangle\langle\Psi_{j}\mid$, where $p_{j}>0$ and $\sum_{j}p_{j}=1$. Entanglement of Formation represents the minimum entanglement resources required to construct this state and reflects that the involving computation is inherently complex. While for bipartite pure-state QE, Entanglement of Formation simplifies to the Partial Von Neumann Entropy (PVNE)
@@
However, for mixed-state QE, in general, this measure heavily relies on the decompositions of pure states, with no general algorithm existing to determine the minimum decomposition to date. To address this issue and improve the Entanglement of Formation behavior in mixed states, Reference wk discovered that for two-qubit mixed states, Entanglement of Formation reduces to
@@
and $C(\rho)$ denoting Concurrence is defined as wk ; au ; ka ; mf ; xjw
@@ II.3 Relative Entropy of Entanglement for Bipartite QE
For bipartite QE especially for bipartite mixed-state QE, another well-known and commonly used bipartite QE measure, Relative Entropy of Entanglement (REE) vv1 ; vv2 was introduced and is defined as
@@
where $\mathcal{S}$ represents a set of separable bipartite density matrix. On one hand, Relative Entropy of Entanglement measures the shortest ”distance” from a given quantum entangled state to the set of separable states, which is the geometric interpretation of this method and reflects that the involving computation is inherently complex. On the other hand, while for bipartite pure QE states, Relative Entropy of Entanglement (REE) simplifies to the Partial Von Neumann Entropy (PVNE)
@@
II.4 Negativity for $2\times 2$ and $2\times 3$ quantum systems
For a disentangled state $\rho$, its partial transpose results in a positive or valid state $\rho^{T_{i}}$, the so-called well-known Peres condition for separable states A.P . For $2\times 2$ and $2\times 3$ quantum states, this condition holds true M.H . It reveals a fact that the partial transpose of a non-separable state $\rho^{T_{i}}$ has one negative eigenvalue $\lambda_{\min}$. Then for $2\times 2$ and $2\times 3$ quantum states, Negativity is defined as twice the absolute value of this negative eigenvalue K.Z ; gv . Mathematically, it is expressed as:
$\operatorname{N}(\rho)=2\max(0,-\lambda_{\min}).$ (10)
@@ II.4 Negativity for $2\times 2$ and $2\times 3$ quantum systems
Summary: Due to the locality of the aforementioned five kinds of bipartite QE measures, symmetry about different partitions is generally observed only for symmetric states. The widely recognized measures for quantifying bipartite mixed-state QE, such as Entanglement of Formation and Relative Entropy of Entanglement, involve determining the minimum QE required for a given state, reflecting the inherent complexity of their computation. While Concurrence is applicable only to $2\times 2$ quantum systems and Negativity limited to $2\times 2$ and $2\times 3$ quantum systems. The organization and structure of bipartite mixed-state QE is more complicated and less well understood than that of pure-state QE. Quantifying QE in mixed states poses additional challenges due to the contributions from classical probability distributions, which complicate both the definition and computation of QE measures. In this context, our research aims to explore new measures for bipartite mixed-state QE in detail, as outlined below.
III Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE)
@@
In any bipartite quantum system, all the information is contained within its wave function or density matrix. When a subsystem is obtained by partial tracing or averaging out the other subsystem freedom, some information is inevitably lost. Therefore, the concept of Reduction-induced Variation of Partial Von Neumann Entropy is presented to quantify bipartite QE, which holds certain importance and practical significance. To elaborate further, let’s consider a bipartite quantum system. In the bipartite density matrix $\rho$, the density matrix of the $i$-th $(i=1,2)$ party, denoted as the bipartite $i$-th party density matrix, only preserves the information of the $i$-th party from $\rho$. This particular density matrix is represented as $\rho(i)$ different from its reduced density matrix $\rho_{i}$. Reduction-induced Variation of Von Neumann Entropy (RIVPVNE) of one-sided density matrix is defined as follows:
$\operatorname{RIVPVNE}(\rho)=S(\rho_{i})-S(\rho(i)),$ (11)
@@
which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to any bipartite states especially for any bipartite mixed states. This measure vanishes if and only if the state is separable, and on average, does not experience an increase under stochastic Local Operations and Classical Communications, thereby establishing itself as an entanglement monotone. Notably, the computational complexity associated with this method is minimal, making it highly efficient. Moreover, it has a wide range of applicability. The intuitive and clear physical picture, combined with simple computation and wide range of applications, facilitates the exploration of its specific potential applications. For any bipartite pure-state QE, this measure reduces to Partial Von Neumann Entropy
@@
Due to the locality of this measure, in general only for symmetric states, it has symmetry about different partition. Subsequently, we provide examples to demonstrate the superiorities of this method in identifying QE through a comparison with other existing bipartite QE measures, such as Partial Von Neumann Entropy, Concurrence, Negativity, Entanglement of Formation and Relative Entropy of Entanglement, considering both their physical implications and mathematical structures, as detailed below.
@@ IV Its Superiorities in comparison to existing bipartite QE measures
with $0\leqslant\alpha\leqslant 2\pi$. The corresponding single-site reduced density matrix and the bipartite single-site density matrix are respectively given by
$\rho_{i}=\left[\begin{array}[]{cc}\cos^{2}\alpha&0\\ 0&\sin^{2}\alpha\\ \end{array}\right]$ (15)
and
$\rho(i)=\left[\begin{array}[]{cc}\cos^{2}\alpha&\cos\alpha\sin\alpha\\ \cos\alpha\sin\alpha&\sin^{2}\alpha\\ \end{array}\right].$ (16)
Consequently, the Reduction-induced Variation of Partial Von Neumann Entropy, as well as the Entanglement of Formation and Relative Entropy of Entanglement, simplifies to the Partial Von Neumann Entropy:
@@
This highlights the interference properties and indicates a clear and intuitive relationship between QE and quantum coherence. The previously mentioned measures are illustrated in Fig.1. For the specified pure states, it is observed that when $\alpha=\frac{\pi}{4},\frac{3\pi}{4}$, all the measures satisfy: $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=1$; whereas for $\alpha=0,\frac{\pi}{2},\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=0$; and for the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{2}<\alpha<\pi$, the relationships $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)<\operatorname{C}(\rho)=\operatorname{N}(\rho)$ hold. Clearly, for the specified pure two-qubit systems, the monotonicities of these measures are fully consistent. The Reduction-induced Variation of Partial Von Neumann Entropy, Entanglement of Formation and Relative Entropy of Entanglement reduce to the Partial Von Neumann Entropy, but exhibit slight differences from Concurrence and Negativity due to variations in their physical implications and mathematical structures.
@@
Fig. 1: Reduction-induced Variation of Partial Von Neumann Entropy as well as other QE measures such as the Partial Von Neumann Entropy, Entanglement of Formation, Relative Entropy of Entanglement, Concurrence and Negativity varying with $\alpha$. When $\alpha=\frac{\pi}{4},\frac{3\pi}{4}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=1$; whereas for the values $\alpha=0,\frac{\pi}{2},\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=0$; and in the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{2}<\alpha<\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)<\operatorname{C}(\rho)=\operatorname{N}(\rho)$.
We next consider the symmetric bipartite mixed states
$\rho=\frac{\sin^{2}\alpha}{2}\mid 10\pm 01\rangle\langle 10\pm 01|+\cos^{2}\alpha|00\rangle\langle 00\mid.$ (19)
@@
The first term corresponds to a QE state with a probability of $\sin^{2}\alpha$, while the second term represents a non-entangled state with a probability of $\cos^{2}\alpha$. Consequently, the bipartite QE is less than $\sin^{2}\alpha$. The corresponding single-site reduced density matrix and bipartite single-site density matrix are
$\rho_{i}=\left[\begin{array}[]{cc}\frac{\sin^{2}\alpha}{2}&0\\ 0&\frac{\sin^{2}\alpha}{2}+\cos^{2}\alpha\\ \end{array}\right]$ (20)
and
$\rho(i)=\left[\begin{array}[]{cc}\frac{\sin^{2}\alpha}{2}&\frac{\sin^{2}\alpha}{2}\\ \frac{\sin^{2}\alpha}{2}&\frac{\sin^{2}\alpha}{2}+\cos^{2}\alpha\\ \end{array}\right].$ (21)
Next, the Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE) is computed as
$\operatorname{RIVPVNE}=S(\rho_{i})-S(\rho(i))$
@@
as well as other QE measures such as the Entanglement of Formation
@@
the Relative Entropy of Entanglement and Concurrence:

### v4

@@ Abstract
The organization and structure of bipartite mixed-state quantum entanglement (QE) are more complex and less well understood compared to bipartite pure-state QE. Bipartite mixed-state QEs and their measures play a crucial role in both theory and practical applications. Some existing measures involve quantifying the minimum QE and reflect the inherently complex nature of their computation, while others are only applicable to highly limited-dimensional quantum systems. In this context, we propose a method termed Reduction-induced Variation of Partial Von Neumann Entropy to quantify QE in any bipartite states, particularly focusing on bipartite mixed states. Partial Von Neumann Entropy is merely a special case of this method, specifically designed to measure QE in bipartite pure states. This method exhibits minimal computational complexity and broad applicability. Its intuitive and clear physical representation, along with easy computation and wide applicability, facilitates exploring its potential applications. Furthermore, we present examples to demonstrate the superiorities of this method in identifying bipartite QE by comparing with other existing bipartite mixed-state QE measures through both their physical implications and mathematical structures.
Keywords:Quantum Entanglement (QE); Bipartite Mixed-state QE; Reduction-induced Variation of Partial Von Neumann Entropy
@@ I Introduction
The type or degree of QEs are often directly related to their revolutionary applications including quantum cryptography kk ; ng , quantum teleportation ch ; rc ; nr , quantum superdense coding nr ; sj , measurement-based quantum computation rr ; fa and quantum enhanced sensing tc ; rcp ; lp . Therefore, a deeper understanding and accurate metric of QE have become important prerequisites for the development of quantum information science and quantum technology. However, to date, the quantum mechanical structure of QE in bipartite mixed states remains unclear; let alone the organization and structure of multipartite QEs. Here, our primary focus is on QEs in bipartite mixed states.
@@
Nonseparable bipartite systems exhibit QE epr ; sd , which can be quantified using various measures. These include the Partial Von Neumann Entropy ch1 ; ch2 , applicable only to bipartite pure states, and Concurrence, which is confined to $2\times 2$ quantum systems wk ; au ; ka ; mf ; xjw . Additionally, Entanglement of Formation ch1 ; ch2 and Relative Entropy of Entanglement vv1 ; vv2 for bipartite quantum states, involve quantifying the minimum QE, and their assessments generally entail considerable computational burdens, especially for bipartite mixed-state QE. In general, QEs in bipartite pure states are more intuitive and easier to comprehend, whereas those in bipartite mixed states are more complicated and less well understood. The quantification of QEs in bipartite mixed states is more complex and challenging due to the contributions from classical probability distributions, making the definition and computation of QEs more intricate.
@@
Mixed states are more reflective of real-world scenarios, as factors like environmental conditions, quantum measurements and thermal influences can lead quantum systems to exist in mixed states. Hence, the quantification of QE in bipartite mixed states is more meaningful in practice, aligning more effectively with experimental conditions and practical applications. Hence, in practical research and applications, bipartite mixed-state QE and its measures hold significant practical importance, in addition to the theoretical significance rooted in the underlying structure of quantum mechanics.
@@
In this context, our research aims to explore new metrics for quantifying bipartite mixed-state QE. We propose a method termed Reduction-induced Variation of Partial Von Neumann Entropy to quantify QE in any bipartite states. Partial Von Neumann Entropy is merely a special case of this method, serving as a measure of bipartite pure-state QE. The computational complexity associated with this approach is considerably low, making it highly efficient. Moreover, it has a wide range of applicability. The intuitive and clear physical representation, along with simple computation and wide range of applications, facilitates the exploration of its specific potential applications. Subsequently, we provide examples to demonstrate the superiorities of this method in bipartite QE identification through a comparison with other established bipartite QE measures such as Partial Von Neumann Entropy, Relative Entropy of Entanglement, Entanglement of Formation and Concurrence considering both their physical implications and mathematical structures.
II existing four kinds of Bipartite QE measures
@@ II.1 Partial Von Neumann Entropy for Bipartite Pure-state QE
where $\rho_{i}(i=1,2)$ represents its subsystem reduced density matrix. This Partial Von Neumann Entropy can effectively measure the degree of QE between bipartite pure states, while for bipartite mixed-state QE, it is ineffective.
@@ II.2 Entanglement of Formation for Bipartite QE and Concurrence for two-qubit QE
For bipartite QE, especially for bipartite mixed-state QE, an extensively acknowledged type of measure, Entanglement of Formation (EOF) ch1 ; ch2 was introduced and is defined as
@@
over all possible ensemble realizations of the bipartite density matrix $\rho=\sum_{j}p_{j}\mid\Psi_{j}\rangle\langle\Psi_{j}\mid$, where $p_{j}>0$ and $\sum_{j}p_{j}=1$. Entanglement of Formation represents the minimum entanglement resources required to construct this state and reflects that the involving computation is inherently complex. While for bipartite pure-state QE, Entanglement of Formation simplifies to the Partial Von Neumann Entropy
@@
However, for mixed-state QE, in general, this measure heavily relies on the decompositions of pure states, with no general algorithm existing to determine the minimum decomposition to date. To address this issue and improve the Entanglement of Formation behavior in mixed states, Reference wk discovered that for two-qubit mixed states, the Entanglement of Formation reduces to
@@
and $C(\rho)$ denotes Concurrence which is defined as wk ; au ; ka ; mf ; xjw
@@ II.3 Relative Entropy of Entanglement for Bipartite QE
Another well-known and commonly used bipartite QE measure, especially for bipartite mixed-state QE, Relative Entropy of Entanglement (REE) vv1 ; vv2 is defined as
@@
where $\mathcal{S}$ represents a set of separable bipartite density matrix. On one hand, Relative Entropy of Entanglement measures the shortest “distance”from a given quantum entangled state to the set of separable states, which is the geometric interpretation of this method and reflects that the involving computation is inherently complex. On the other hand, while for bipartite pure QE states, the Relative Entropy of Entanglement simplifies to the Partial Von Neumann Entropy
@@
The four kinds of bipartite QE measures mentioned above exhibit various flaws and limitations. For example, Partial Von Neumann Entropy is limited to pure bipartite quantum systems; Entanglement of Formation and Relative Entropy of Entanglement involve determining the minimum QE required for a given state, reflecting the inherent complexity of their computation. While Concurrence is applicable only to $2\times 2$ quantum systems. The organization and structure of bipartite mixed-state QE are more complicated and less well understood than those of pure-state QE. Quantifying QE in mixed states poses additional challenges due to the contributions from classical probability distributions, which complicate both the definition and computation of QE measures. In this context, our research aims to explore new measures for bipartite mixed-state QE in detail, as outlined below.
III Reduction-induced Variation of Partial Von Neumann Entropy
@@
In any bipartite QE system, all the information is contained within its wave function or density matrix. When a subsystem is obtained by partial tracing or averaging out the freedom of the other subsystem, some information is inevitably lost. Therefore, the concept of Reduction-induced Variation of Partial Von Neumann Entropy is presented to quantify bipartite QE, which holds specific importance and practical significance. In the bipartite density matrix $\rho$, the density matrix of the $i$-th $(i=1,2)$ party, only preserves all the information of the $i$-th party from $\rho$; this bipartite $i$-th part density matrix is denoted as $\rho(i)$ and differs from its reduced density matrix $\rho_{i}$. To elaborate further, let’s begin with bipartite qubit systems. As is well known, the Bell entangled states are
$\Psi_{\pm}=\frac{\sqrt{2}}{2}|11\rangle\pm\frac{\sqrt{2}}{2}|00\rangle,$ (10)
and
$\Phi_{\pm}=\frac{\sqrt{2}}{2}|10\rangle\pm\frac{\sqrt{2}}{2}|01\rangle.$ (11)
And their corresponding density matrices are respectively
$\rho=\left[\begin{array}[]{cccc}\frac{1}{2}&0&0&\frac{1}{2}\\ 0&0&0&0\\ 0&0&0&0\\ \frac{1}{2}&0&0&\frac{1}{2}\\ \end{array}\right]$ (12)
with the corresponding bipartite single-site density matrix and the single-site reduced density matrix provided as follows:
$\rho(i)=\left[\begin{array}[]{cc}\frac{1}{2}&\frac{1}{2}\\ \frac{1}{2}&\frac{1}{2}\\ \end{array}\right]$ (13)
and
$\rho_{i}=\left[\begin{array}[]{cc}\frac{1}{2}&0\\ 0&\frac{1}{2}\\ \end{array}\right];$ (14)
and
$\rho=\left[\begin{array}[]{cccc}0&0&0&0\\ 0&\frac{1}{2}&\frac{1}{2}&0\\ 0&\frac{1}{2}&\frac{1}{2}&0\\ 0&0&0&0\\ \end{array}\right]$ (15)
with the corresponding bipartite single-site density matrix and the single-site reduced density matrix specified as follows:
$\rho(i)=\left[\begin{array}[]{cc}\frac{1}{2}&\frac{1}{2}\\ \frac{1}{2}&\frac{1}{2}\\ \end{array}\right]$ (16)
and
$\rho_{i}=\left[\begin{array}[]{cc}\frac{1}{2}&0\\ 0&\frac{1}{2}\\ \end{array}\right].$ (17)
For the Bell entangled states, the Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE) is expressed as:
$\operatorname{RIVPVNE}(\rho)=S(\rho_{i})-S(\rho(i))=S(\rho_{i}).$ (18)
Inspired by this, we consider a general entangled bipartite qubit system. The density matrix that best reflects the QE of the bipartite qubit must satisfy the following form,
$\rho_{\operatorname{eff}}=\left[\begin{array}[]{cccc}\rho_{11,11}&0&0&\rho_{11,00}\\ 0&\rho_{10,10}&\rho_{10,01}&0\\ 0&\rho_{01,10}&\rho_{01,01}&0\\ \rho_{00,11}&0&0&\rho_{00,00}\\ \end{array}\right],$ (19)
that is to say
$\rho_{11,10}=\rho_{10,11}=\rho_{11,01}=\rho_{01,11}$
$=$ $\rho_{00,10}=\rho_{10,00}=\rho_{00,01}=\rho_{01,00}=0,$ (20)
which serves as a necessary condition. The system’s density matrix can take different forms based on different single-site orthonormal basis vectors $\{\cos\theta|1\rangle-\sin\theta|0\rangle,\cos\theta|0\rangle+\sin\theta|1\rangle,0\leqslant\theta<2\pi\}$}. Therefore, we first select an appropriate orthonormal basis such that the density matrix is expressed in a form that most effectively represents the entangled state $\rho_{\operatorname{eff}}$. Under these conditions, the entropy of the bipartite $i$-th part density matrix reaches its maximum value $S_{\operatorname{Max}}(\rho(i)_{\operatorname{eff}})$ depending on the choice of single-site orthonormal basis vectors. The same approach is also applied to bipartite high-dimensional quantum systems. The Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE) of one-sided density matrix is defined as follows:
$\operatorname{RIVPVNE}(\rho)=S(\rho_{i})-S_{\operatorname{Max}}(\rho(i)_{\operatorname{eff}}),$ (21)
@@
This measure vanishes if and only if the state is separable, and on average, does not experience an increase under Stochastic Local Operations and Classical Communications, thereby establishing itself as an entanglement monotone G.V ; Y.C . Consequently, it can be employed to quantify the QE degree in bipartite systems. Notably, the computational complexity associated with this method is minimal, making it highly efficient. Moreover, it has a wide range of applicability. The intuitive and clear physical picture, combined with simple computation and wide range of applications, facilitates the exploration of its specific potential applications. For any bipartite pure-state QE, this measure reduces to the Partial Von Neumann Entropy
@@
Partial Von Neumann Entropy is merely a special case of this method, serving as a measure of bipartite pure-state QE. Due to the locality of this measure, in general only for symmetric states, it has symmetry about different partition. Subsequently, we provide examples to demonstrate the superiorities of this method in identifying QE through a comparison with other existing bipartite QE measures such as Partial Von Neumann Entropy, Relative Entropy of Entanglement, Entanglement of Formation and Concurrence taking into account both their physical implications and mathematical structures, as detailed below.
@@ IV Its Superiorities in comparison to existing bipartite QE measures
with $0\leqslant\alpha\leqslant\frac{\pi}{2}$. Consequently, the Reduction-induced Variation of Partial Von Neumann Entropy, as well as the Entanglement of Formation and Relative Entropy of Entanglement, simplifies to the Partial Von Neumann Entropy:
@@
This highlights the interference properties and indicates a clear and intuitive relationship between QE and quantum coherence. The previously mentioned measures or witnesses are illustrated in Fig.1. For the specified pure states, it is observed that when $\alpha=\frac{\pi}{4}$, all the measures satisfy: $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=1$; whereas for $\alpha=0,\frac{\pi}{2}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0$; and for the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{4}<\alpha<\frac{\pi}{2}$, the relationships $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)<\operatorname{C}(\rho)$ hold. Clearly, the monotonicities of these measures are fully consistent. The Reduction-induced Variation of Partial Von Neumann Entropy, Entanglement of Formation and Relative Entropy of Entanglement all reduce to the Partial Von Neumann Entropy, but is less than or equal to Concurrence which here signifies quantum coherence, because of differences in their physical implications and mathematical structures.
@@
Fig. 1: Reduction-induced Variation of Partial Von Neumann Entropy as well as other QE measures such as the Partial Von Neumann Entropy, Entanglement of Formation, Relative Entropy of Entanglement and Concurrence varying with $\alpha$. When $\alpha=\frac{\pi}{4}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=1$; whereas for the values $\alpha=0,\frac{\pi}{2}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0$; and in the ranges $0<\alpha<\frac{\pi}{4}$ and $\frac{\pi}{4}<\alpha<\frac{\pi}{2}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)<\operatorname{C}(\rho)$.
We next consider the symmetric bipartite states
$\rho=\frac{\sin^{2}\alpha}{2}\mid 10\pm 01\rangle\langle 10\pm 01|+\cos^{2}\alpha|00\rangle\langle 00\mid$ (27)
@@
with $0\leqslant\alpha\leqslant\pi$. The first term corresponds to a pure-state QE with a probability of $\sin^{2}\alpha$, while the second term represents a non-entangled pure state with a probability of $\cos^{2}\alpha$. Hence, the bipartite mixed-state QE is less than $\sin^{2}\alpha$. Next, the Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE) is computed as
$\operatorname{RIVPVNE}=S(\rho_{i})-S_{\operatorname{Max}}(\rho(i)_{\operatorname{eff}})$
@@
in addition to other QE measures like the Entanglement of Formation given by
@@
and the Relative Entropy of Entanglement as well as Concurrence given by